When I add 3 times my age 3 years from now to 4 times my age 4 years from now, I get 8 times my current age. How old will I be 2 years from now
Answer by stanbon(75887)

When I add 3 times my age 3 years from now to 4 times my age 4 years from now, I get 8 times my current age. How old will I be 2 years from now
-----
Let your age be "x".
Equation:
3(x+3) + 4(x+4) = 8x
3x+9 + 4x+16 = 8x
x = 25
----
x+2 = 27 yrs (your age in 2 years)
=============================
Cheers,
Stan H.
=============================